OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate 1/0.502 as follows: Let f(x)=1/x and find the equation of the tangent line to f(x) at a "nice" point near 0.502. Then use this to approximate 1/0.502.

OpenStudy (anonymous):

so a tangent line is parallel at the point (0.502), and inorder to find the tangent line we need the slope, m. m = (y2-y1) / (x2-x1) so lets pick 2 points near our point of 0.502, or inother words, lets do linear approximation. so solving 1/x where x1 = 0.501 and then do x2 = 0.503 now we get two points (0.501, 1.9960) (0.503, 1.9881) can you find the slope from here with these two points then find the equation of the tangent line?

OpenStudy (anonymous):

I got -3.95 as the slop

OpenStudy (anonymous):

yes, now use the y=mx+b formula. you now have m, but now you need x and y, so use the x point given you, 0.502 and solve for y so, 1/0.502 = y. now you have m, x, and y, now you can solve for b in the y=mx+b equation

OpenStudy (anonymous):

What do I do once I get b?

OpenStudy (anonymous):

rewrite the equation of y=mx+b, and replace m and b with the values you found

OpenStudy (anonymous):

so, f(x) = -3.95x -1.0046

OpenStudy (anonymous):

Okay, I got that

OpenStudy (anonymous):

is there anything else you need help with?

OpenStudy (anonymous):

I'm not sure where to go after that...